A medical practitioner wants to estimate the mean systolic blood pressure in pre-school children with inbom heart disease. He plans on using a 99% confidence interval and a margin of error of 5 units. The standard deviation of systolic blood pressures in pre-school children with inborn heart disease is 20. How many children should be enrolled in the study?


1
Expert's answer
2022-06-09T14:35:31-0400

"E=z_{0.005}\frac{\sigma}{\sqrt{n}}"

"n=(\frac{z_{0.005}\sigma}{E})^2=(\frac{2.576*20}{5})^2=107"